# DCT-Shield: A Robust Frequency Domain Defense against Malicious Image Editing

### v1

@@ 5.4 Implementation Details
We run DCT-Shield with the following default settings. Immunization quality is set to $Q_{alg}=0.95$ and bound adversarial perturbation in the quantized DCT-coefficients to $\epsilon=1$ across all three channels. The step size $\gamma=0.1$ The algorithm is run for 1000 iterations. We work with images of resolution $512\times 512$. We clearly state different settings in the text in case we change any of the settings for an ablation study.
@@ 6.1 Immunization Quality
We assess DCT-Shield’s immunization quality by evaluating (i) the perceptibility of added adversarial noise and (ii) its effectiveness in preventing malicious edits. For comparison, baseline methods use a standard pixel budget ($L_{\infty}$-norm) of $16/255$ pixels, while DCT-Shield operates at default settings ($Q_{alg}=0.95$, $\epsilon=1$) as described in Sec. 5.4. Unlike baselines, DCT-Shield does not impose an explicit pixel budget, but this is not a limitation, as RGB space norms do not reliably reflect human visual perception [18, 25]. Thus, pixel budgets are not a suitable metric for noise perceptibility.
Qualitative results. Figure 3 showcases qualitative comparisons across various editing tasks. DCT-Shield introduces significantly less perceptible adversarial noise while providing strong protection against edits, including environment changes (first row), object replacement (second row), and object addition (third row). Notably, edits on DCT-Shield’s immunized images fail to preserve key semantics, ensuring effective protection, whereas baselines sometimes fall short. Additional examples, including attribute and style modifications, are in Section C of the supplementary material.
Quantitative results. We also show quantitative results in Table 1 to concretely showcase DCT-Shield’s immunization quality. DCT-Shield typically outperforms baseline methods across various metrics, in terms of both edit-protection and noise imperceptibility. Notably, in cases where DCT-Shield does not provide the best perception levels, it indeed does so in the corresponding protection metric and vice versa. This indicates a trade-off between the two aspects and is discussed in the next paragraph. Additionally, human evaluation scores indicate a clear preference for DCT-Shield’s immunized images and its ability to safeguard against edits. Details of human evaluation studies are presented in Section D2 of the supplementary material.
Protection vs. noise-imperceptibility trade-off. Comparing perturbations in DCT coefficients to those in pixel space is challenging. To ensure fairness, we evaluate DCT-Shield at different coefficient perturbation limits $\epsilon=[0.6,0.8,1,1.2,1.4]$ and and baselines at pixel-space budgets of $1,2,4,8$ and $16$ ($/255$). Fig. 5 illustrates the trade-off between edit protection and noise perceptibility across immunization methods. Noise perceptibility is measured by the FID score between input and immunized images (lower is better), while edit protection is evaluated by the FID score between edited versions (higher is better). DCT-Shield achieves a superior balance, providing strong edit protection with minimal perceptible noise.
@@ 6.2 Immunization Against Malicious Inpainting
Defense against inpainting attacks poses a significant challenge, as an attacker can mask out a sensitive region from an image and seamlessly blend it into a new background. Effective protection requires adding targeted perturbations to these sensitive regions to disrupt inpainting models. To mitigate this threat, we employ DCT-Shield with a quality parameter of $Q_{alg}=0.9$, applying masked noise exclusively to the Y channel. This configuration ensures strong inpainting protection while preserving the imperceptibility of the added noise. All baseline methods are run with a pixel budget of $16$. We evaluate our approach against baseline methods in Table 2, where results demonstrate that DCT-Shield outperforms or is at least comparable across all evaluation metrics. A qualitative comparison of inpainting results is presented in Figure 4, illustrating that our method provides stronger protection by effectively preventing the attacker’s intent. Notably, we achieve these results using only encoder-based attacks, whereas several baselines rely on diffusion-based attacks, which are computationally more intensive. Additional qualitative results and an analysis of JPEG robustness for this task are provided in Section C of the supplementary material.
@@ 6.3 Robustness to Purification Techniques
Adversarial noise purification techniques aim to remove immunization perturbations while preserving data integrity. Common methods include JPEG compression and crop & resize, while advanced approaches like AdvClean [^1] https://github.com/lllyasviel/AdverseCleaner use image filtering. Malicious editors may exploit these techniques to reverse immunization. This section evaluates DCT-Shield’s robustness against such purification methods, comparing it to baseline approaches.
JPEG Conversion. We evaluate edit protection after JPEG purification at quality levels 65, 75, 85, and 95 percent. Fig. 6 shows qualitative results, where DCT-Shield (only-Y) at $Q_{alg}=0.85$. We show qualitative results in Fig. 6. demonstrates significantly higher robustness than baseline methods. The figure highlights DCT-Shield’s superior capability of preserving semantics of the edit of its immunized image, and is evident from the semantics of the cat in the last four rows of Fig. 6.
Quantitative results in Fig. 7 confirm DCT-Shield’s resilience using LPIPS, with similar trends observed in other metrics. Unlike pixel-space methods, DCT-Shield integrates JPEG pipeline in noise optimization, ensuring that images immunized generated with $Q_{alg}=q$ remain largely unaffected by JPEG purification at quality levels exceeding $q$. Furthermore, this enables defenders to fine-tune $Q_{alg}$ to optimize the trade-off between robustness range and noise perceptibility. Notably, such adaptability is not attainable through pixel-space optimization methods. Qualitative results showing range robustness are shown in Section C of the supplementary material
Crop and Resize. Following the approach in [9, 4], immunized images were cropped by 64 pixels at the boundaries and then resized to their original dimensions $512\times 512$. The edits of these cropped and resized immunized images were compared to the edits of the original input. As shown in Fig. 8, DCT-Shield demonstrates greater robustness to this crop-and-resize purification method.
@@ 7 Conclusion
In this work, we introduced DCT-Shield, a novel adversarial immunization technique that leverages the Discrete Cosine Transform (DCT) domain to add imperceptible yet effective noise for protecting images against malicious edits. By leveraging the JPEG pipeline in the optimization process, DCT-Shield achieves robustness against many purification methods while maintaining high visual fidelity, and offers adaptability through JPEG-aware and JPEG-averse immunization. Furthermore, we presented multiple DCT-Shield variants, including Mask-based and Y-channel DCT-Shield, tailored for the inpainting task. Extensive qualitative and quantitative evaluations demonstrate that DCT-Shield outperforms prior methods in both edit protection and imperceptibility. DCT-Shield employs significantly fewer parameters than pixel-space approaches. Furthermore, by utilizing only a VAE in the optimization process, it remains computationally efficient and independent of specific U-Net architectures across different editing models. This makes DCT-Shield a practical and scalable solution for real-world image security. Future research can explore extending DCT-Shield to video data and enhancing its robustness against emerging purification techniques.

### v2

@@ 5.4 Implementation Details
We run DCT-Shield with the following default settings. Immunization quality is set to $Q_{alg}=0.95$ and the adversarial perturbation in the quantized DCT-coefficients is bound by $\epsilon=1$ across all channels. The step size $\gamma=0.1$ The algorithm is run for 1000 iterations. We work with images of resolution $\times 512$ 512. We clearly state different settings in the text if they are changed for an ablation or experiment.
@@ 6.1 Immunization Quality
We assess DCT-Shield’s immunization quality by evaluating (i) the perceptibility of added adversarial noise and (ii) its effectiveness in preventing malicious edits. For comparison, baseline methods use a standard pixel budget ($L_{\infty}$-norm) of 16/25516/25516 / 255 pixels, while DCT-Shield operates at default settings ($Q_{alg}=0.95$, $\epsilon=1$) as described in Sec. 5.4. Unlike baselines, DCT-Shield does not impose an explicit pixel budget, but this is not a limitation, as RGB space norms do not reliably reflect human visual perception [19, 26]. Thus, pixel budgets are not a suitable metric for noise perceptibility.
Qualitative results. Figure 3 showcases qualitative comparisons across various editing tasks. DCT-Shield introduces significantly less perceptible adversarial noise while providing strong protection against edits, including environment changes (first row), object replacement (second row), and object addition (third row). Notably, edits on DCT-Shield’s immunized images fail to preserve key semantics, ensuring effective protection, whereas baselines sometimes fall short. More qualitative results across edit-types are shown in Section D of the supplementary material.
Quantitative results. We also show quantitative results in Table 1 to concretely showcase DCT-Shield’s immunization quality. DCT-Shield typically outperforms baseline methods across various metrics, in terms of both edit-protection and noise imperceptibility. Notably, in cases where DCT-Shield does not provide the best perception levels, it indeed does so in the corresponding protection metric and vice versa. This indicates a trade-off between the two aspects and is discussed in the next paragraph. Additionally, human evaluation scores indicate a clear preference for DCT-Shield’s immunized images and its ability to safeguard against edits. Details of human evaluation studies are presented in Section C.2 of the supplementary material.
Protection vs. noise-imperceptibility trade-off. Comparing perturbations in DCT coefficients to those in pixel space is challenging. To ensure fairness, we evaluate DCT-Shield at different coefficient perturbation limits $\epsilon=[0.8,1,1.2,1.4]$ and baselines at pixel-space budgets of 10,12,1410,12,1410 , 12 , 14 and 161616 (/255/255/ 255). Fig. 5 illustrates the trade-off between edit protection and noise perceptibility across immunization methods. Noise perceptibility is measured by the FID score between input and immunized images (lower is better), while edit protection is evaluated by the FID score between edited versions (higher is better). DCT-Shield achieves a superior balance, providing strong edit protection with minimal perceptible noise.
@@ 6.2 Immunization Against Malicious Inpainting
Defense against inpainting attacks poses a significant challenge, as an attacker can mask out a sensitive region from an image and seamlessly blend it into a new background. Effective protection requires adding targeted perturbations to these sensitive regions to disrupt inpainting models. To mitigate this threat, we employ DCT-Shield with a quality parameter of $Q_{alg}=0.9$, applying masked noise to the YCbCr channels. This configuration ensures strong inpainting protection while preserving the imperceptibility of the added noise. All baseline methods are run with a pixel budget of 161616. We evaluate our approach against baseline methods in Table 2, where results demonstrate that DCT-Shield outperforms or is at least comparable across all evaluation metrics. A qualitative comparison of inpainting results is presented in Figure 4, illustrating that our method provides stronger protection by effectively preventing the attacker’s intent. Notably, we achieve these results using only encoder-based optimization, whereas several baselines rely on diffusion-based optimization, which is computationally more intensive. Additional qualitative results for this task are shown in Section D of the supplementary material.
@@ 6.3 Robustness to Purification Techniques
Adversarial noise purification techniques aim to remove immunization perturbations while preserving data integrity. Common methods include JPEG compression and crop & resize, while advanced approaches like AdvClean [^2] https://github.com/lllyasviel/AdverseCleaner use image filtering. Malicious editors may exploit these techniques to reverse immunization. This section evaluates DCT-Shield’s robustness against such purification methods, comparing it to baseline approaches. Results for purification methods like Gaussian noising, Impress[3] and noisy upscaling are shown in Section D.3 of the supplementary.
JPEG Conversion. We evaluate edit protection after JPEG purification at quality levels 65, 75, 85, and 95 percent. Fig. 6 shows qualitative results with DCT-Shield (only-Y) at $Q_{alg}=0.85$. The figure highlights DCT-Shield’s superior capability of retaining image immunity despite JPEG conversions. This is evident from the semantics of the cat in the last column (DCT-Shield) vs. other columns (baselines) across all JPEG quality values.
Quantitative results in Fig. 7 confirm DCT-Shield’s resilience using LPIPS, with similar trends observed in other metrics. Unlike pixel-space methods, DCT-Shield integrates JPEG pipeline in noise optimization, ensuring that images immunized generated with $Q_{alg}=q$ remain largely unaffected by JPEG purification at quality levels exceeding $q$. Furthermore, this enables defenders to fine-tune $Q_{alg}$ to optimize the trade-off between robustness range and noise perceptibility. Notably, such adaptability is not attainable through pixel-space optimization methods. Qualitative results showing range robustness are shown in Section D.4 of the supplementary material.
Crop and Resize. Following the approach in [10, 5], immunized images were cropped by 64 pixels at the boundaries and then resized to their original dimensions $\times 512$ 512. The edits of these cropped and resized immunized images were compared to the edits of the original input. As shown in Fig. 8, DCT-Shield demonstrates greater robustness to this crop-and-resize purification method.
@@ 7 Conclusion
In this work, we present DCT-Shield - a novel image immunization technique that operates in the Discrete Cosine Transform (DCT) domain to embed imperceptible yet effective perturbations, protecting images from malicious edits. By leveraging the JPEG pipeline during optimization, DCT-Shield achieves robustness against a wide range of purification methods while preserving high visual fidelity. It also allows users to control the balance between protection-robustness and noise-imperceptibility through adjustable parameters. We further introduced multiple DCT-Shield variants, designed to address task-specific requirements and improve parameter efficiency. Extensive qualitative and quantitative evaluations demonstrate that DCT-Shield outperforms prior methods in both edit protection and imperceptibility. DCT-Shield employs significantly fewer parameters than pixel-space approaches. Furthermore, by utilizing only a VAE in the optimization process, it remains computationally efficient and independent of specific U-Net architectures across different editing models. This makes DCT-Shield a practical and scalable solution for real-world image security. Future research can explore extending DCT-Shield to video data and enhancing its robustness against emerging purification techniques.